the oldest child in the family of four children is three times as old as the youngest. the two middle children are 21 and 22 years old. if the average age of the children is 27.75, how old is the youngest child?

Answer by TimothyLamb(4379)

---
x = youngest
y = mid younger
z = mid older
u = oldest
---
u = 3x
y = 21
z = 22
(x + y + z + u)/4 = 27.75
---
put the system of linear equations into standard form:
---
3x - u = 0
y = 21
z = 22
x + y + z + u = 111
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = youngest = 17
y = mid younger = 21
z = mid older = 22
u = oldest = 51
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php